10 POI I Calculate the future worth of a $1,000 (invested today) 6 years from now if the rate of return is 10%.
Answer #1

Method 1: compound interest formula

Future value formula

Present value= 1000

N= 6 years

Rate of return = 10%

FV= PV (1+r)^n

FV= 1000(1+0.10)^6

FV= 1000(1.1)^6= 1000× 1.771561

FV= 1771.56

Method 2:

As per simple interest

I=P r*n

1000× 0.10×6= 600

Then future worth =principle + interest = 1000+600= 1600